The width of a rectangle is
fixed at 8 ft. What lengths will make the perimeter at
least 200 ft? at most 200 ft?
My text does not really explain how to work this problem. could you please demonstrate how to solve this?

Answer by mananth(16949)

width = 8 feet
let length be x feet
..
Perimeter = at least 200 .Meaning, It cannot be less than 200.
2(L+8)>=200
/2
L+8 >= 200/2
L+8>=100
-8
l>=100-8
l>=92 feet
...
Similarly at most 200 means : maximum 200. it can be less than 200
2(l+8)<=200
I am sure you will walk through now.
..
m.ananth@hotmail.ca